Consider a glass of 160 mL of water at 27°C. Calculate the mass of ice at -15°C that must be added to cool the water to 10°C after thermal equilibrium is achieved. To find the mass of water use the density of water = 1.0 g/mL.
_____ g
Answer #1

Let a gram ice is added, and let heat capacity of water = S j/g.°C

Change in temperature for water = 27 - 10 = 17

Mass of 160 ml water = 160 gram

Heat lost by water = 160S× 17 = 2720S

Change in temperature for ice